kekeman:

Which of the following tables represents a linear function? x 2 2 2 2 2 y −3 −2 −1 0 1 x −2 −1 0 2 4 y −4 −2 −1 0 1 x −6 −2 0 1 3 y −7 negative one third −5 negative four thirds 1 x −3 −1 1 3 5 y −7 negative nine halves −2 one half 3

silvernight269:

Find the slope or you can just pair them into coordinates and see if they become linear

KyledaGreat:

Let's go through each table and see if it represents a linear function. For a table to represent a linear function, there must be a constant rate of change between any two points. We can check this by looking at the differences in $y$ values when we move from one $x$ value to the next. If the differences are constant, then the function is linear. Let's start with the first table. x | 2 | 2 | 2 | 2 | 2 --|---|---|---|---|-- y |-3 |-2 |-1 |0 |1 When we move from $x=2$ to $x=2$ (the same $x$ value), the $y$ value changes from $-3$ to $-2$, so the rate of change is $1$. However, when we move from $x=2$ to $x=2$ again, the $y$ value changes from $-2$ to $-1$, so the rate of change is $1$ again. This means that the function is not linear, since the rate of change is not constant. Let's move on to the second table. x |-2 |-1 |0 |2 |4 --|---|---|---|---|-- y |-4 |-2 |-1 |0 |1 When we move from $x=-2$ to $x=-1$, the $y$ value changes from $-4$ to $-2$, so the rate of change is $2$. When we move from $x=-1$ to $x=0$, the $y$ value changes from $-2$ to $-1$, so the rate of change is $1$. When we move from $x=0$ to $x=2$, the $y$ value changes from $-1$ to $0$, so the rate of change is $1/2$. When we move from $x=2$ to $x=4$, the $y$ value changes from $0$ to $1$, so the rate of change is $1/2$ again. Since the rate of change is not constant, the function is not linear. Let's move on to the third table. x |-6 |-2 |0 |1 |3 --|---|---|---|---|-- y |-7 |-1/3 |-5 |-4/3 |1 When we move from $x=-6$ to $x=-2$, the $y$ value changes from $-7$ to $-1/3$, so the rate of change is $22/3$. When we move from $x=-2$ to $x=0$, the $y$ value changes from $-1/3$ to $-5$, so the rate of change is $-5/2$. When we move from $x=0$ to $x=1$, the $y$ value changes from $-5$ to $-4/3$, so the rate of change is $7/3$. When we move from $x=1$ to $x=3$, the $y$ value changes from $-4/3$ to $1$, so the rate of change is $13/3$. Since the rate of change is not constant, the function is not linear. Finally, let's look at the fourth table. x |-3 |-1 |1 |3 |5 --|---|---|---|---|-- y |-7 |-9/2 |-2 |1/2 |3 When we move from $x=-3$ to $x=-1$, the $y$ value changes from $-7$ to $-9/2$, so the rate of change is $5/2$. When we move from $x=-1$ to $x=1$, the $y$ value changes from $-9/2$ to $-2$, so the rate of change is $5/2$ again. When we move from $x=1$ to $x=3$, the $y$ value changes from $-2$ to $1/2$, so the rate of change is $7/4$. Since the rate of change is not constant, the function is not linear. Therefore, none of the tables represents a linear function.

surjithayer:

let the equation of linear function be y=ax+b 1.it is not a function as x=2 gives different values of y 2.x=0,y=-1 -1=a*0+b b=-1 y=ax-1 x=2,y=0 0=2a-1 2a=1 a=1/2 y=1/2 x-1 x=4,y=1 1=1/2*4-1 1=2-1 1=1 (true) x=-2,y=-4 -4=1/2*(-2)-1 -4=-1-1 -4=-1-1 -4=-2 false it is not a linear function. 3. y=ax+b x=0,y=-5 -5=a*0+b b=-5 y=ax-5 x=-6 y=-7 -7=-6a-5 6a=-5+7 6a=2 a=2/6=1/3 y=1/3 x-5 x=3,y=1 1=1/3*3-5 1=1-5 1=-4 false not a linear function. 4. x=-3 y=-7 -7=-3a+b .... (-1) x=5,y=3 3=5a+b ....(2) subtract(1) from (2) 10=8a a=10/8=5/4 from (1) -7=-3(5/4)+b b=-7+15/4=(-28+15)/4=-13/4 y=5/4 x-13/4 x=3,y=1/2 1/2=5/4(3)-13/4 1/2=15/4-13/4=2/4=1/2 true x=-1,=-9/2 -9/2=(5/4)*(-1)-13/4=-5/4-13/4=-18/4=-9/2 true x=1,y=-2 -2=5/4(1)-13/4=-8/4=-2 true Hence fourth is alinear function.

KyledaGreat:

Yes, your analysis is correct. The fourth equation is a linear function, while the first three are not linear functions. You have correctly found the slope and y-intercept of the fourth equation, and have tested it with three different points to confirm that it is a linear function. Well done!